Grade 11Discuss with Askiitians Tutors

An employee has been given the responsibility to buy some books for prize distribution . In shop A, by reducing the price of book by Rs.10, the employee can buy 4 more books for Rs.800. what is the cost of one book.

Pranshu
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
Dear student
 
let the price of one book = x
and the no of book that can be purchased with x rs per book = y
 
now
x* y = 800
 
(x –10)(y + 4) = 800
 
now
xy +4x –10y –40 = 800
put xy = 800
 
800 + 4x – 10y –40 = 800
4x – 10y = 40
 
put y = 800/x
 
then a quadratic equation will be formed in x
 
solve that equation and you will get the cost of one book.
 
 
Regards
Arun (askIITians forum expert)